OpenStudy (khally92):

Hello am new here need ur help.. For the function f(x)=2xcube-7xsquare+4x+1 a) Find the instanteneous rate of change at x=0 and x=1 b) Is the function increasing or decreasing at x=0? how about at x=1? c) Do you expect a local miximum or local minimum when 0

OpenStudy (khally92):

yeah i got stock with x=0

OpenStudy (turingtest):

\[f(x)=2x^3-7x^2+4x+1\]and that is not the formula for instantaneous rate of change above

OpenStudy (turingtest):

instantaneous rate of change is found through taking the derivative

OpenStudy (henryblah):

Just find f'(x) and sub in the x values

OpenStudy (anonymous):

Oh. That's true. Sorry. That's the average rate of change.

OpenStudy (turingtest):

so @khally92 what did you get for the derivative?

OpenStudy (khally92):

6xsquare-14x+4

OpenStudy (turingtest):

and so what did you get for f'(1) and f'(0) ???

OpenStudy (khally92):

f(1)= -4 and f(0)=4

OpenStudy (turingtest):

and those are the answers for a) nice job :) to answer part b) we just need to remember that f(x) is decreasing is when f'<0 f(x) is increasing is when f'>0 so what is the answer to that part?

OpenStudy (khally92):

f>0 so that means its increase and at x=1 is decreasing

OpenStudy (khally92):

at -4 is decreasing and at 4 is increasing.. hello

OpenStudy (turingtest):

sorry at x=1, f'<0 so it's decreasing yes

OpenStudy (turingtest):

an yes to the other question too now what do you think about part c)

OpenStudy (khally92):

no graph to conclude whether 0<x<1 is a local mini or local maxi

OpenStudy (henryblah):

Put f'(x)=0 Find the turning point within the range. Then find the 2nd derivative and sub that value in. If <0 then max, if >0 min.

OpenStudy (henryblah):

You can use your answer to part b. If it is increasing first then decreasing, then max. If it is the other way round then min.

OpenStudy (khally92):

its increasing decreasing so its a local max ryt?

OpenStudy (henryblah):

Yes, if that is the case

OpenStudy (khally92):

yes... thank you thank you very much n @turningpoint too thnks alot

OpenStudy (henryblah):

TuringTest=Turningpoint :D

OpenStudy (khally92):

another one please differentiate f(x)= squareroot of x-square of 3. can u please write it out let me see if its correct

OpenStudy (anonymous):

Is this what the equation looks like x^(3)^(1/2)^(1/2)? ^(1/2) is the same as squaring. Just use exponent rules and simplify to get (1/2)*(1/2)*3 = 3/4 so x^(3/4) then differentiate that as you normally would using nx^n-1. Know what I mean?

OpenStudy (khally92):

yeah i know that power rule is that what am going to use.. this is the question again squareroot of x - squareroot sign3outside then x inside hope u gog the equation ryt.. i got 1/2suare root of x-1/3square root of x is tha correct

OpenStudy (anonymous):

can you use numbers and ^ for to the power I cannot understand what you are trying to show

OpenStudy (khally92):

well here is the equation xraisedpower of 1/2 - xraised power of -1/3... hope that helps

OpenStudy (anonymous):

Yeah I can understand that a little better it just isnt a very good way of writing the question. You just use power rule. Do you know how to subtract from a fraction? Post your answer I will tell you if it is correct.

OpenStudy (khally92):

x raisedpower of -1/2 divide by2 - xraised power of -2/3 divide by 3. which i can futher simplyfy

OpenStudy (khally92):

thank you...

OpenStudy (anonymous):

No problem you solved it.

OpenStudy (khally92):

Last for the day the mass mass of M in gramms of undissolved sugar left in a tea cup after t seconds is given by M=10.5-o.4tsquare. when will the sugar dissolve?

OpenStudy (anonymous):

You are looking for the time when mass is equal to zero so what do you think you should set your mass to in your equation?

OpenStudy (khally92):

looking fo time when mass is zero...so i should find a value for t that will make m zero?e m zero?

OpenStudy (anonymous):

Yes and inorder to find the t that make m=0 you just set M to 0 in the equation and then solve for t.

OpenStudy (khally92):

thats squareroot of 26.25

OpenStudy (anonymous):

Correct.

OpenStudy (khally92):

Back to differenciation 3x/x2+4... am going to use the product rule ryt when x2+4comes up having the negative sign could u please help me differenciate the (x2+4)-1..

OpenStudy (anonymous):

okay easy you will also need chain rule. Just do 3x*(x^2+4)^-1 like you said and use product rule to get 3x'*(x^2+4)^-1+3x*(x^2+4)^-1' try it yourself and I will help you from there.

OpenStudy (khally92):

well thats (3x)(2x)^-1 + (3)(x^2+4)

OpenStudy (anonymous):

Okay lets back up.

OpenStudy (khally92):

well thats (3x)(2x)^-1 + (3)(x^2+4)^-1

OpenStudy (anonymous):

Okay it seems like you understand the product rule but not the chain rule. for (x^2+4)^-1 first you take the derivatie of the outside then you multiply that by the derivative of the inside understand?

OpenStudy (khally92):

thats-1x

OpenStudy (anonymous):

Okay here I will help you (-1)(x^2+4)^-2*(2x) do you understand how I did this?

OpenStudy (khally92):

yeah i do kind of look loke caine rule

OpenStudy (anonymous):

It is the chain rule haha.

OpenStudy (anonymous):

you want to show the whole problem solved to see if you got it or do you feel confident you have it?

OpenStudy (khally92):

Find the equation of the tangent to y=x2-3x-4 is parrelto the line y=7x+3..... i jus need to know if the answer i got is correct. i have y=7x-29 points on the curve arex=5 and y=6 are the answers correct. good evening

OpenStudy (anonymous):

yes you got it